# A Study of Undefined Behavior Across Foreign Function Boundaries in Rust Libraries

### v5

@@ Abstract
Developers rely on the static safety guarantees of the Rust programming language to write secure and performant applications. However, Rust is frequently used to interoperate with other languages which allow design patterns that conflict with Rust’s evolving aliasing models. Miri is currently the only dynamic analysis tool that can validate applications against these models, but it does not support foreign functions, indicating that there may be a critical correctness gap across the Rust ecosystem. We conducted a large-scale evaluation of Rust libraries that call foreign functions to determine whether Miri’s dynamic analyses remain useful in this context. We used Miri and an LLVM interpreter to jointly execute applications that call foreign functions, where we found 47 instances of undefined or undesired behavior from 37 libraries. Three bugs were found in libraries that had more than 10,000 daily downloads on average during our observation period, and one was found in a library maintained by the Rust Project. Many of these bugs were violations of Rust’s aliasing models, but the latest Tree Borrows model was significantly more permissive than the earlier Stacked Borrows model. The Rust community must invest in new, production-ready tooling for multi-language applications to ensure that developers can detect these errors.
@@
Rust, C, C++, interoperation, undefined behavior, aliasing, bugs, foreign functionsI Introduction
@@
Miri [3] is a widely-used Rust interpreter that uses dynamic analysis to detect violations of Rust’s aliasing model. A limitation of Miri is that it cannot detect these errors across foreign function boundaries. However, one of the most common reasons for using Rust’s unsafe features is to call foreign functions [4, 5, 6, 7, 8]. We seek to determine whether the differences between Rust and other languages can lead to errors in practice.
@@
To answer these questions, we created MiriLLI: a tool which combines Miri with an LLVM interpreter to jointly execute programs and detect undefined behavior across foreign function boundaries. We used MiriLLI to conduct a large-scale study of 9,130 test cases from 957 Rust libraries that call foreign functions. We identified 47 instances of undefined or undesirable behavior in 37 of these libraries. Of the 90 test cases that violated Stacked Borrows, 66% (59) did not violate Tree Borrows.
@@
In Section II, we compare Rust’s semantics with C and C++, and we describe the resources and best practices that Rust developers use to interoperate with these languages. In Section III, we document our methodology for sampling and evaluating test cases from Rust libraries that call foreign functions, and we describe the challenges that we encountered when implementing MiriLLI. In Section IV, we describe each type of bug that we found. We discuss the implications of our findings in Section V. We review prior work on Rust interoperation in Section VI, we discuss threats to validity in Section VII, and we conclude in Section VIII. Our dataset, the Appendix, and the source for MiriLLI are available in our replication package111https://github.com/icmccorm/mirilli.
@@ II-A Rust’s Aliasing Model
Rust developers can avoid these issues by ensuring that their programs adhere to the rules of Rust’s aliasing model. The Tree Borrows [10] model provides the latest definition of these rules. Under this model, each pointer has provenance that determines its permission to read or write to a location [15]. Each location is associated with a tree that tracks all valid permissions to its contents. When a location is borrowed, a branch is created in the tree that holds a new permission to access the location. When a pointer is used, its permission is checked against the tree to ensure that it is valid and that it permits the kind of access that takes place. Each type of access will have a different effect on the tree depending on the position of each permission in the tree relative to the permission used for the access. With respect to a given permission, an access is considered a child access if it requires that permission or any of its descendants; otherwise, it is a foreign access. Accessing a location using an invalid permission is considered undefined behavior.
We created three categories to classify the Tree Borrows violations that we observed in our investigation. Figure II-A provides minimal examples of each one. We refer to the first category as an expired permission error. Under Tree Borrows, writing to a memory location may cause other permissions to that location to expire (or transition to , in terms of the model). Using one of these expired permissions is undefined behavior. We provide a minimal example of this in Figure 2, where the variable x is mutably borrowed twice; once by y and again by z. However, we also store a copy of y as a raw pointer (*mut). Rust’s borrow checker does not place aliasing restrictions on raw pointers, so it assumes that z has exclusive, mutable access from when it is created on line 4 until it is used on line 6. Writing through y on line 5 breaks this assumption, so writing through z is now undefined behavior. In terms of the model, the permissions for y and z are children of the permission for x. Both receive a permission, which has the capability to become a unique permission for writing to x. When a permission is used for a write access, it becomes an permission, and all foreign permissions become , canceling their reservations. Since z is adjacent to y, the write on line 5 is a foreign access.
@@
The third type of undefined behavior that we observed is a framing error. We provide an example in Figure 1(c), where the function free receives a mutable reference. This reference encodes the promise that ownership will eventually be returned to the caller. However, this promise is broken when the reference is cast into a raw pointer and used to free the underlying allocation. In terms of Tree Borrows, when a reference-type value is passed as an argument to a function, it becomes protected, making it undefined behavior for it to transition to ---even if it is never accessed after this point. When a function returns, the permissions of its arguments are no longer protected.
@@ II-B1 Initialization
Rust’s borrow checker prevents variables from being used until they are initialized, and no primitive types tolerate indeterminate values in safe contexts. However, Rust’s standard library provides the MaybeUninit<T> struct, which represents an instance of T that may or may not be initialized. Once an instance has been fully initialized, the unsafe function MaybeUninit<T>::assume_init() will unwrap the outer struct to produce the value inside. It is undefined behavior to call assume_init() if T is not fully initialized.
@@ II-B3 Provenance
Both C and C++ allow pointers to be converted to and from integers with a size equal to the word size of the current architecture.555C23 §6.3.2.3.5-6; C++23 §6.7.1.10.4-5 Implementations may track the provenance, or origin of pointers to inform compile-time optimizations [19], but neither of the current standards for C or C++ defines if or how provenance should be preserved across these conversions. Rust’s pointers have provenance [15], but the specifics of Rust’s provenance model are undecided.
Miri has multiple methods for handling provenance across pointer-to-integer conversion. By default, when a pointer is converted to an integer, its tag is added to a set of tags that have been ‘‘exposed’’ by pointer-to-integer conversion. When integers are converted back into pointers, their allocation identifier is reconstructed from their address, but they receive a wildcard provenance value. Under Stacked Borrows, when a memory access occurs using a wildcard tag, Miri eagerly interprets it as being equivalent to any tag in the stack that permits the access. The current implementation of Tree Borrows does not support this behavior yet; all accesses with wildcard tags are allowed and do not affect the state of the tree. Alternatively, developers can enable strict provenance, which treats integer-to-pointer conversion as an error.
@@ II-C Interoperation
Developers can access foreign functions and static variables using Rust’s foreign function interface. Declarations, or ‘‘bindings,’’ for each of these objects are written in
extern blocks, which take an optional qualifier to indicate the Application Binary Interface (ABI) used by the foreign library. Developers can also use the
extern keyword to change the ABI of a function defined in Rust so that it can be called from other languages. Only a subset of Rust’s types are guaranteed to be compatible with foreign ABIs. Both structs and enums must be annotated with the
#[repr(C)] attribute to ensure that their layout is compatible with C. The Rust compiler enforces the use of ABI-compatible types via lints, but these can be disabled. Bindings are not validated against their definitions, but tools such as bindgen [20] and CXX [21] can automatically generate bindings from headers.
@@ III-A Sampling
We also used this opportunity to partially measure the distribution of crates that interact with foreign code. When we compiled each crate, we used an early lint pass to examine the abstract syntax tree and record the locations of all function declarations with a foreign ABI. We used data from this pass to identify 2,516 crates that declared a foreign function. This represents only 2.1% of all valid crates. However, 12,564 (10.4%) of all valid crates had a version that directly depended on at least one prior version of a crate that declared a foreign function binding. The maximum number of dependents for any crate that declared foreign functions was 6,411, with a mean of 11.0 dependents and a standard deviation of 171.8 dependents.
@@ III-B Implementation
When MiriLLI encounters a foreign function call in a Rust program, it looks for a corresponding definition in the LLVM module. After finding a definition, the function’s arguments are passed through a translation function, which lowers them into the representation used by LLI. There are two translation functions indicated by the arrows in the middle of Figure 2; one from Miri’s representation to LLI’s (left) and another from LLI’s representation to Miri’s (right). Miri passes the translated arguments to LLI, which creates a new thread to execute the function. Then, Miri sets the current current Rust thread to join on this new LLVM thread. After the LLVM thread terminates, its return value is passed back through the translation layer and given to the Rust thread, which continues to execute. A similar process is used when an LLVM thread calls a Rust function. Although Miri is single-threaded, it supports concurrency by non-deterministically stepping through multiple simulated ‘‘threads’’ of execution. LLI did not originally support any form of multithreading, but we modified it to be compatible with Miri’s implementation. We support calling foreign functions from parallel Rust threads, but we do not support multithreading within LLVM.
@@ III-C Evaluation
After deduplication, we had 394 errors to investigate. Multiple factors lead to the number of bugs being significantly smaller than the number of deduplicated test outcomes. Our method of deduplication was conservative, we ignored errors from crates that had been unpublished since the start of our investigation, and we only reported Stacked Borrows violations that were also Tree Borrows violations. We also observed false positives related to alignment and accesses through addresses that we do not emulate (e.g., stdout).
@@
Not all errors are undefined behavior, and not all instances of undefined behavior are readily exploitable. However, we still attempted to follow ethical vulnerability disclosure practices by reporting bugs privately via email before creating public reports. When we found a bug, we examined its crate’s Cargo.toml and the GitHub profile of its repository’s owner to find a way to contact its maintainers privately. If we were unsuccessful, we logged a public issue. We also logged issues if we had not received a response after one month or if the type of error did not appear to be exploitable. Our reports included a representative test case, the output from MiriLLI, and a minimal example where applicable. When a fix was trivial, we also filed a pull request. Section 1 of our Appendix includes metadata for each library and links to our contributions.
@@ IV Results
We found 47 new bugs from 37 crates. As of July 29th, 2024, 22 of the bugs have been fixed, and none have been identified as security vulnerabilities. Bugs occurred slightly more frequently in LLVM than in Rust; 25 occurred in LLVM, 16 occurred in Rust, and 6 were caused by incorrect Rust bindings to foreign functions. The majority (34) were found in crates with less than 100 average daily downloads in the six months prior to our snapshot of crates.io. However, we discovered three Tree Borrows violations in separate crates that each had more than 10,000 average daily downloads. We have no reason to believe that any of these bugs would have been found by an unmodified version of Miri. We found several additional errors that occurred in Rust and did not meaningfully involve the FFI. These errors are excluded from the counts here, but appear in our raw dataset.
We found 90 tests from 37 crates where Stacked Borrows violations occurred in foreign code. However, 66% (59) of these tests passed or encountered an unsupported operation under Tree Borrows. This is mainly because 50% (45) had invalid range errors similar to the minimal example shown earlier in Figure II-A. We found several instances where slices of an array were cast into raw pointers, passed across the FFI, and then offset beyond the range that they initially borrowed. This is considered undefined behavior under Stacked Borrows, but it is valid under Tree Borrows. This provides our answer to RQ2; Tree Borrows accepts more real-world programs that call foreign functions than Stacked Borrows due to differences in handling pointer arithmetic.
@@
| LLVM | LLVM | 1 | 3 | - | 4 |
| Rust | LLVM | 1 | 16 | 1 | 18 |
| Rust | Rust | 10 | 2 | 4 | 16 |
| Total: | 12 | 24 | 11 | 47 | |
@@ IV-A Ownership
We found 24 Ownership bugs from 20 crates. These include 18 Tree Borrows violations and six accesses out-of-bounds.
@@
We found 10 insufficient permission errors caused by incorrectly casting immutable references to mutable raw pointers. In Bug #0, a chain of casts was used to convert &self into *mut T. This pointer had a read-only permission, so mutating this structure across the FFI triggered an insufficient permission error. There were six bugs with equivalent errors. The four remaining bugs involved foreign function bindings that incorrectly declared pointers as const instead of mut, which led developers to pass pointers derived from immutable references to these functions instead of mutable ones.
@@
In Rust, it is not possible to create a struct that contains a self-reference using only safe code. Allowing this could lead to use-after-free errors depending on the order in which each field is dropped. It is possible, but non-trivial, to implement self-referential patterns using safe encapsulations of unsafe operations [31]. Bug #0 demonstrates how a self-referential pattern can be implemented incorrectly. We found this bug in an encapsulation of a C library implementing a minimal file system for embedded applications. It used separate structs to represent the state of the file system and its configuration. The configuration held a mutable reference to the state, but both objects were contained in a parent struct representing the entire file system. Rust held a single mutable reference to the parent, which became invalid when the state was mutated through the configuration by a foreign function.
Figure IV-A provides a minimal example of the bug and our fix. Lines we removed are highlighted in and marked with ‘‘-’’, while lines we added are highlighted in and marked with a ‘‘+’’. Tags within a circle indicate the state of the tree prior to the fix, while tags within a square represent the state of the tree after the fix. When a tag is on a line, it indicates that the pointer or reference on that line has a permission within the subtree corresponding to the tag. The tables at the bottom display the state of the tree before and after the write access on line 18.
The Rust encapsulation for open initializes the field buffer of the struct Alloc on line 10 with a mutable raw pointer cast from a mutable reference to the field cache. The struct is mutably borrowed again on line 11, creating a reference that is passed into the foreign function ffi::open. Each of these borrows, indicated by and , corresponds to an adjacent branch in the tree. Across the FFI, a write access occurs on line 18 using . This is a child access relative to and , so they transition to . However, it is foreign access relative to , so this permission becomes . This leads to an expired permission error on line 14 when is used for a read access in Rust.
We fixed this bug by wrapping parts of the state in an UnsafeCell, which can be mutated through shared references. In Figure IV-A, we change the first field of Alloc to UnsafeCell<i32> and replace the mutable borrow assigned to a.buffer with a.cache.get(). This expression does not perform a retag, so both a and cache share permission . Since a contains an UnsafeCell, the mutable borrow assigned to b on line 11 associates the tag with a special permission that can tolerate foreign writes. Consequently, this permission is no longer disabled by the write through on line 18, so the read access on line 14 is valid.
@@
Expired permission errors occurred when pointers derived from mutable references were copied into the foreign heap. We observed this type of bug in five separate libraries. The most notable examples were Bug #0 from bzip2 and Bug #0 from flate2, which are both popular compression libraries. In particular, flate2 is actively maintained by the Rust project and averaged more than 130,000 downloads per day during our observation period.
Figure IV-A shows a minimal example of the cyclic aliasing pattern used by each library. We use the same notation as Figure IV-A. The structs Stream and State carry pointers to each other. Both are encapsulated in Rust by a Compression object, which carries a heap-allocated instance of Stream. The Rust function new allocates memory for a Stream object and passes a mutable reference carrying the permission to ffi::init. This foreign function copies the pointer it receives to the Stream object into a newly allocated State instance on line 22, creating a cyclic structure. Later, when the function compress is called, it creates an adjacent permission . Mutating through this permission on line 15 causes to become , making it undefined behavior to read using in the body of ffi:compress on line 25.
To fix each error, we unwrapped the Box using into_raw() and stored a raw pointer to the allocation within the Compression struct. We modified the API so that this raw pointer would be used for every access to the allocation, ensuring that the cyclic structure would remain valid. This corresponds to permission in Figure IV-A, which remains through each access. We also had to modify Compression’s implementation of Drop to rewrap this pointer using Box::into_raw() to avoid a memory leak.
@@
Framing
We found one framing violation equivalent to the example in Figure 1(c). In Bug #0, a Rust function attempted to use a foreign function to grow a heap-allocated array of values. The function was passed a pointer derived from a reference of type &’static mut [T], which was protected. Similar to Figure 1(c), when this pointer was used to reallocate the array, its protected permission became , which is undefined behavior. We fixed this bug by changing the function to receive *mut [T]. Raw pointers do not receive protectors, so they can be used for deallocation.
@@
Rust’s PhantomData<T> is a ‘‘ghost type;’’ it behaves like an instance of T, but it can be created without providing a concrete value of that type. However, there is an exception: a value of type PhantomData<UnsafeCell<T>> does not behave like an UnsafeCell<T>; it cannot be mutated through immutable references [32]. In Bug #0, the type Opaque was declared as an alias for PhantomData<UnsafeCell<* mut T>>. It was used to represent foreign types in Rust without having to declare their layout on both sides of the FFI boundary. Since Opaque did not contain a concrete instance of UnsafeCell, raw pointers derived from &Opaque received a read-only permission, leading to an insufficient permission error when they were used for a write access in foreign code. We proposed fixing this error by removing PhantomData, but other contributors are still discussing the correct fix. The relationship between PhantomData and UnsafeCell is not yet settled, so this pattern may be allowed in the future.
@@
We found six accesses out-of-bounds in unique crates. Bugs #0 and #0 were found in two encapsulations for libdecnumber, which is part of the GCC toolchain. This library often used the following pattern to iterate over arrays of integers:
@@
Bug #0 involved a foreign function that expected to receive a pointer to a string with six characters. It was allocated with CString::new(Vec::with_capacity(6)), exposed as a raw pointer using into_raw(), and correctly deallocated after the call using from_raw. However, the implementation of CString::new appends a null-terminator to the vector and then converts it into a value of type Box<[u8]> using into_boxed_slice(), which shortens the capacity of a vector to be equal to its length. The intended capacity was six, but into_boxed_slice() shortened it to one, appending a null-terminator. The encapsulation did not validate that the length of the CString instance was correct, leading to an access out-of-bounds when the C codebase read beyond the first character. We fixed this error by initializing the CString with a constant of the correct length.

### v6

@@ Abstract
Developers rely on the static safety guarantees of the Rust programming language to write secure and performant applications. However, Rust is frequently used to interoperate with other languages which allow design patterns that conflict with Rust’s evolving aliasing models. Miri is currently the only dynamic analysis tool that can validate applications against these models, but it does not support finding bugs in foreign functions, indicating that there may be a critical correctness gap across the Rust ecosystem. We conducted a large-scale evaluation of Rust libraries that call foreign functions to determine whether Miri’s dynamic analyses remain useful in this context. We used Miri and an LLVM interpreter to jointly execute applications that call foreign functions, where we found 46 instances of undefined or undesired behavior in 37 libraries. Three bugs were found in libraries that had more than 10,000 daily downloads on average during our observation period, and one was found in a library maintained by the Rust Project. Many of these bugs were violations of Rust’s aliasing models, but the latest Tree Borrows model was significantly more permissive than the earlier Stacked Borrows model. The Rust community must invest in new, production-ready tooling for multi-language applications to ensure that developers can detect these errors.
@@
Rust, interoperation, undefined behavior, aliasing, bugs, foreign functionsI Introduction
@@
Miri [3] is a widely-used Rust interpreter that uses dynamic analysis to detect violations of Rust’s aliasing model. A limitation of Miri is that it cannot detect these errors across foreign function boundaries. However, one of the most common reasons for using Rust’s unsafe features is to call foreign functions [4, 5, 6, 7, 8]. We seek to determine whether the differences between Rust and other languages are leading to errors in practice.
@@
To answer these questions, we created MiriLLI: a tool which combines Miri with an LLVM interpreter to jointly execute programs and detect undefined behavior across foreign function boundaries. We used MiriLLI to conduct a large-scale study of 9,130 test cases from 957 Rust libraries that call foreign functions. We identified 46 unique instances of undefined or undesirable behavior from 37 libraries. Of the 90 test cases that violated Stacked Borrows, 66% (59) did not violate Tree Borrows.
@@
In Section II, we compare Rust’s semantics with C and C++, and we describe the resources and best practices that Rust developers use to interoperate with these languages. In Section III, we document our methodology for sampling and evaluating test cases from Rust libraries that call foreign functions, and we describe the challenges that we encountered when implementing MiriLLI. In Section IV, we describe each type of bug that we found. We discuss the implications of our findings in Section V. We review prior work on Rust interoperation in Section VI, we discuss threats to validity in Section VII, and we conclude in Section VIII. Our dataset, the Appendix, and the source for MiriLLI are available in our replication package111https://doi.org/10.5281/zenodo.12727039.
@@ II-A Rust’s Aliasing Model
Rust developers can avoid these issues by ensuring that their programs adhere to the rules of Rust’s aliasing model. The Tree Borrows [10] model provides the latest definition of these rules. Under this model, each pointer has provenance that determines its permission to read or write to a location [15]. Each location is associated with a tree that tracks all valid permissions to its contents. When a location is borrowed, a branch is created in the tree that holds a new permission to access the location. When a pointer is used, its permission is checked against the tree to ensure that it is valid and that it permits the kind of access that takes place. Each type of access will have a different effect on the tree depending on the position of each permission in the tree relative to the permission used for the access. With respect to a given permission, an access is considered a child access if it requires that permission or any of its descendants; otherwise, it is a foreign access. Accessing a location using an invalid permission is undefined behavior.
There are three categories of aliasing violations under Tree Borrows that we observed in our investigation. Figure II-A provides minimal examples of each one. We refer to the first category as an expired permission error. Under Tree Borrows, writing to a memory location may cause other permissions to that location to expire (or transition to , in terms of the model). Using one of these expired permissions is undefined behavior. We provide a minimal example of this in Figure 2, where the variable x is mutably borrowed twice; once by y and again by z. However, we also store a copy of y as a raw pointer (*mut). Rust’s borrow checker does not place aliasing restrictions on raw pointers, so it assumes that z has exclusive, mutable access from when it is created on line 4 until it is used on line 6. Writing through y on line 5 breaks this assumption, so writing through z is now undefined behavior. In terms of the model, the permissions for y and z are children of the permission for x. Both receive a permission, which has the capability to become a unique permission for writing to x. When a permission is used for a write access, it becomes an permission, and all foreign permissions become , canceling their reservations. Since z is adjacent to y, the write on line 5 is a foreign access.
@@
The third type of undefined behavior that we observed is a protected permission error. We provide an example in Figure 1(c), where the function free receives a mutable reference. This reference encodes the promise that ownership will eventually be returned to the caller. However, this promise is broken when the reference is cast into a raw pointer and used to free the underlying allocation. In terms of Tree Borrows, when a reference-type value is passed as an argument to a function, it becomes protected, making it undefined behavior for it to transition to ---even if it is never accessed after this point. When a function returns, the permissions of its arguments are no longer protected.
@@ II-B1 Initialization
Rust’s borrow checker prevents variables from being used until they are initialized, and no primitive types tolerate indeterminate values in safe contexts. However, Rust’s standard library provides MaybeUninit<T>, which represents an instance of T that may or may not be initialized. Once an instance has been fully initialized, the unsafe function MaybeUninit<T>::assume_init will unwrap the outer struct to produce the value inside. It is undefined behavior to call assume_init if T is not fully initialized.
@@ II-B3 Provenance
Both C and C++ allow pointers to be converted to and from integers with a size equal to the word size of the current architecture.555C23 §6.3.2.3.5-6; C++23 §6.7.1.10.4-5 Implementations may track the provenance, or origin of pointers to inform compile-time optimizations [19]. However, neither of the current standards for C or C++ defines if or how provenance should be preserved across these conversions, though several models have been proposed [20]. Rust’s pointers have provenance [15], but the specifics of Rust’s provenance model are undecided.
Miri has multiple methods for handling provenance across pointer-to-integer conversion. By default, when a pointer is converted to an integer, its tag is added to a set of tags that have been ‘‘exposed’’ by pointer-to-integer conversion. When integers are converted back into pointers, their allocation identifier is reconstructed from their address, but they receive a ‘‘wildcard’’ provenance value [21]. Under Stacked Borrows, when a memory access occurs using a wildcard tag, Miri eagerly interprets it as being equivalent to any tag in the stack that permits the access. The current implementation of Tree Borrows does not support this behavior yet; all accesses with wildcard tags are allowed and do not affect the state of the tree. Alternatively, developers can enable strict provenance, which treats integer-to-pointer conversion as an error.
@@ II-C Interoperation
Developers can access foreign functions and static variables using Rust’s foreign function interface. Declarations, or ‘‘bindings,’’ for each of these objects are written in extern blocks, which take an optional qualifier to indicate the Application Binary Interface (ABI) used by the foreign library. Developers can also use the extern keyword to change the ABI of a function defined in Rust so that it can be called from other languages. Only a subset of Rust’s types are guaranteed to be compatible with foreign ABIs. Both structs and enums must be annotated with the #[repr(C)] attribute to ensure that their layout is compatible with C. Bindings are not validated against their definitions, but tools such as bindgen [22] and CXX [23] can automatically generate bindings from headers.
@@ III-A Sampling
We also partially measured the distribution of crates that interact with foreign code. When we compiled each crate, we used an early lint pass to examine the abstract syntax tree and record the locations of all function declarations with a foreign ABI. We used data from this pass to identify 2,516 crates that declared a foreign function. This represents only 2.1% of all valid crates. However, 12,564 (10.4%) of all valid crates had a version that directly depended on at least one prior version of a crate that declared a foreign function binding. The maximum number of dependents for any crate that declared foreign functions was 6,411, with a mean of 11.0 dependents and a standard deviation of 171.8 dependents.
@@ III-B Implementation
When MiriLLI encounters a foreign function call in a Rust program, it looks for a corresponding definition in the LLVM module. After finding a definition, the function’s arguments are passed through a translation function, which lowers them into the representation used by LLI. There are two translation functions indicated by the arrows in the middle of Figure 2; one from Miri’s representation to LLI’s (left) and another from LLI’s representation to Miri’s (right). Miri passes the translated arguments to LLI, which creates a new thread to execute the function. Then, Miri sets the current current Rust thread to join on this new LLVM thread. After the LLVM thread terminates, its return value is passed back through the translation layer and given to the Rust thread, which continues to execute. A similar process is used when an LLVM thread calls a Rust function. Although Miri is single-threaded, it supports concurrency by non-deterministically stepping through multiple simulated ‘‘threads’’ of execution. LLI did not originally support any form of multithreading, but we modified it to be compatible with Miri’s implementation. We support calling foreign functions from parallel Rust threads, but we do not support multithreading within LLVM. We implemented the Rust interface to LLI as an extension of Inkwell: a high-level Rust encapsulation for LLVM [29].
@@ III-C Evaluation
After deduplication, we had 394 errors to investigate. Multiple factors lead to the number of bugs being significantly smaller than the number of deduplicated test outcomes. Our method of deduplication was conservative, we ignored errors from crates that had been unpublished since the start of our investigation, and we only reported Stacked Borrows violations that were also Tree Borrows violations. We also observed false positives related to alignment and accesses through addresses that we do not emulate (e.g., stdout). Additional manual investigation was necessary to fully diagnose certain aliasing bugs (e.g. #0), since MiriLLI currently does not record the source locations of valid updates to the state of the borrow tracker when they occur in LLVM. Miri typically provides a trace of the history of each permission involved in a aliasing violation; only partial traces were available for certain bugs.
@@
Not all errors are undefined behavior, and not all instances of undefined behavior are readily exploitable. However, we still attempted to follow ethical vulnerability disclosure practices by reporting bugs privately via email before creating public reports. When we found a bug, we examined its crate’s Cargo.toml and the GitHub profile of its repository’s owner to find a way to contact its maintainers privately. If we were unsuccessful, we logged a public issue. We also logged issues if we had not received a response after at least one month or if the type of error did not appear to be exploitable. Our reports typically included a representative test case, the output from MiriLLI, and a minimal example where applicable. When a fix was trivial, we also filed a pull request. Section 1 of our Appendix includes metadata for each library and links to our contributions.
@@ IV Results
We found 46 bugs in 37 crates. All appear to be new. As of December 13th, 2024, 28 of the bugs have been confirmed and fixed, and none have been identified as security vulnerabilities. Bugs occurred slightly more frequently in LLVM than in Rust; 23 occurred in LLVM, 17 occurred in Rust, and six were related to incorrect bindings. The majority (32) were found in crates with less than 100 average daily downloads in the six months prior to our snapshot of crates.io. However, we discovered three Tree Borrows violations in separate crates that each had more than 10,000 average daily downloads. We found several additional errors that occurred in Rust and did not meaningfully involve foreign code. These errors are excluded from the counts here, but they appear in our raw dataset.
We found 90 tests from 37 crates where Stacked Borrows violations occurred in foreign code. However, 66% (59) of these tests passed or encountered an unsupported operation under Tree Borrows. This is mainly because 50% (45) had invalid range errors similar to the minimal example shown earlier in Figure II-A. For example, we observed that slices of an array would be cast into raw pointers, passed across the FFI, and then offset beyond the range that they initially borrowed. This is considered undefined behavior under Stacked Borrows, but it is valid under Tree Borrows. This provides our answer to RQ2; Tree Borrows accepts more real-world programs that call foreign functions than Stacked Borrows due to differences in handling pointer arithmetic.
@@
| LLVM | LLVM | 3 | - | 3 | |
| Rust | LLVM | 1 | 16 | - | 17 |
| Rust | Rust | 9 | 2 | 6 | 17 |
| Total: | 10 | 24 | 12 | 46 | |
@@ IV-A Ownership
We found 24 Ownership bugs from 20 crates. These include 17 Tree Borrows violations and seven accesses out-of-bounds.
@@
We found 10 insufficient permission errors caused by incorrectly casting immutable references to mutable raw pointers. In Bug #0, a chain of casts was used to convert &self into a mutable raw pointer. This pointer still had read-only permission, so mutating through it across the FFI triggered an insufficient permission error. There were seven bugs with equivalent errors. The remaining three bugs occurred due to foreign function bindings that incorrectly declared pointers as const instead of mut, leading developers to pass pointers derived from immutable references instead of mutable ones.
@@
In Rust, it is possible, but nontrivial, to implement self-referential patterns using safe encapsulations of unsafe operations [34]. Bug #0 demonstrates how a self-referential pattern can be implemented incorrectly. We found this bug in an encapsulation of a C library implementing a minimal file system for embedded applications. It used separate structs to represent the state of the file system and its configuration. The configuration held a mutable reference to the state, but both objects were contained in a parent struct representing the entire file system. Rust held a single mutable reference to the parent, which became invalid when the state was mutated through the configuration by a foreign function.
Figure IV-A provides a minimal example of the bug and our fix. Lines we removed are highlighted in and marked with ‘‘-’’, while lines we added are highlighted in and marked with a ‘‘+’’. Tags within a circle indicate the state of the tree prior to the fix, while tags within a square represent the state of the tree after the fix. When a tag is on a line, it indicates that the pointer or reference on that line has a permission within the subtree corresponding to the tag. The tables at the bottom display the state of the tree before and after the write access on line 17.
The Rust encapsulation for open initializes the field buffer of the struct Alloc on line 10 with a mutable raw pointer cast from a mutable reference to the field cache. The struct is mutably borrowed again on line 11. Each of these borrows, indicated by and , corresponds to an adjacent branch in the tree. The reference assigned to b, which has permission , is used to access the field buffer, which has permission . The pointer in this field is passed into the foreign function ffi::open. Across the FFI, a write access occurs on line 17 using permission . This is a child access relative to and , so they transition to . However, it is foreign access relative to , so this permission becomes . This leads to an expired permission error on line 14 when is used for a read access in Rust.
We fixed this bug by wrapping parts of the state in an UnsafeCell, which can be mutated through shared references. In Figure IV-A, we change the first field of Alloc to UnsafeCell<i32> and replace the mutable borrow assigned to a.buffer with a.cache.get(). This expression does not perform a retag, so both a and cache share permission . Since a contains an UnsafeCell, the mutable borrow assigned to b on line 11 associates the tag with a special permission that can tolerate foreign writes. Consequently, this permission is no longer disabled by the write through on line 17, so the read access on line 14 is valid.
@@
Expired permission errors occurred when pointers derived from mutable references were copied into the foreign heap. We found this type of bug in five separate libraries. The most notable examples were Bug #0 from bzip2 and Bug #0 from flate2, which are both popular compression libraries. In particular, flate2 is actively maintained by the Rust project and averaged more than 130,000 downloads per day during our observation period.
Figure IV-A shows a minimal example of the cyclic aliasing pattern used by each library. We use the same notation as Figure IV-A. The structs Stream and State carry pointers to each other. Both are encapsulated in Rust by a Compression object, which carries a heap-allocated instance of Stream within a Box. The Rust function new allocates memory for a Stream object and passes a mutable reference carrying the permission to ffi::init. This foreign function copies the pointer it receives to the Stream object into a newly allocated State instance on line 22, creating a cyclic structure. Later, when the function compress is called, it creates an adjacent permission . Mutating through this permission on line 15 causes to become , making it undefined behavior to read using in the body of ffi:compress on line 25.
To fix each error, we unwrapped the Box using into_raw and stored a raw pointer to the allocation within the Compression struct. We modified the API so that this raw pointer would be used for every access to the allocation, ensuring that the cyclic structure would remain valid. This corresponds to permission in Figure IV-A, which remains through each access. We also had to modify Compression’s implementation of Drop to rewrap this pointer using Box::into_raw to avoid a memory leak.
@@
Protection
We found one protected permission error equivalent to the example in Figure 1(c). In Bug #0, a Rust function attempted to use a foreign function to grow a heap-allocated array of values. The foreign function was passed a pointer derived from a reference-type argument to the Rust encapsulation. Similar to Figure 1(c), when this pointer was used to reallocate the array, its protected permission became , which is undefined behavior. We fixed this bug by changing the function to receive a raw pointer instead. Raw pointers do not receive protectors, so they can be used for deallocation.
@@
Rust’s PhantomData<T> behaves like an instance of T, but it can be created without providing a concrete value. However, a value of type PhantomData<UnsafeCell<T>> does not behave like an UnsafeCell<T>; it cannot be mutated through immutable references [35]. In Bug #0, the type Opaque was declared as an alias for PhantomData<UnsafeCell<* mut T>>. It was used to represent foreign types in Rust without having to declare their layout on both sides of the FFI boundary. Since Opaque did not contain a concrete instance of UnsafeCell, raw pointers derived from &Opaque received a read-only permission, leading to an insufficient permission error when they were used for a write access in foreign code. We proposed fixing this error by removing PhantomData, but other contributors are still discussing the correct fix. The relationship between PhantomData and UnsafeCell is not yet settled, so this pattern may be allowed in the future.
@@
We found seven access out-of-bounds errors in unique crates. Bugs #0 and #0 were found in two encapsulations for libdecnumber, which is part of the GCC toolchain [36]. This library often used the following pattern to iterate over arrays of integers:
@@
Bug #0 involved a foreign function that expected to receive a pointer to a string with six characters. It was allocated with CString::new(Vec::with_capacity(6)), exposed as a raw pointer using into_raw, and correctly deallocated after the call using from_raw. However, the implementation of CString::new appends a null-terminator to the vector and then converts it into a value of type Box<[u8]> using into_boxed_slice, which shortens the capacity of a vector to be equal to its length. The intended capacity was six, but into_boxed_slice shortened it to one, appending a null-terminator. The encapsulation did not validate that the length of the CString instance was correct, leading to an access out-of-bounds when the C codebase read beyond the first character. We fixed this error by initializing the CString with a constant of the correct length.